# Nonparametric extensions of nuclear equations of state: probing the breakdown scale of relativistic mean-field theory

### v1

@@ I Introduction
Uncovering the equation of state (EoS) of dense matter is central to understanding neutron stars, bridging nuclear physics and astrophysics Oppenheimer and Volkoff (1939); Rhoades and Ruffini (1974); Lattimer and Prakash (2001); Lovato et al. (2022); Chatziioannou et al. (2024). While the EoS is in principle determined entirely by the standard model, in practice, there is currently no perturbative or exact approach to compute the properties of matter in the densest regions of neutron star cores. Therefore, for the purposes of interpreting macroscopic astronomical observables such as neutron star masses and radii, “model-agnostic” representations of the EoS minimize phenomenological model-dependence Landry and Essick (2019); Essick et al. (2020a); Legred et al. (2022). However, fully-agnostic approaches can overestimate uncertainties in regions where calculations from models of nuclear matter are reliable, for example, near nuclear saturation density ($n_{0}$), where a description of matter with nucleon degrees of freedom is applicable. In addition, extracting detailed microphysical information from model-agnostic approaches is not straightforward, because it is not usually clear what exact microscopic interaction is responsible for determining the EoS at different densities. Certain quantities are not accessible at all, such as the single-particle energies required to compute transport properties Schmitt and Shternin (2018). In contrast with model-agnostic strategies, a variety of approaches have been used to model dense matter up to very high densities, but these come with varying degrees of (often difficult to quantify) systematic uncertainty Detar and Kunihiro (1989); Jido et al. (2001); Buballa (2005); Bender et al. (2003); Dutra et al. (2014).
Chiral effective field theory Weinberg (1990) ($\chi$EFT) provides constraints on neutron-rich matter that are systematically improvable, but calculations cannot be extended to the cores of astrophysical neutron stars Tews et al. (2013); Drischler et al. (2020a); Tews et al. (2018). $\chi$EFT’s nature as an effective theory allows us to systematically write down a series of interaction terms and compute an associated truncation uncertainty. However, its validity is limited to densities $\lesssim 2\,n_{0}$, which does not extend to neutron star cores. Relativistic mean-field theories (RMFTs), in contrast, are usable at densities and temperatures relevant in neutron stars and their mergers Walecka (1974); Boguta and Bodmer (1977); Serot and Walecka (1986); Glendenning (1996); Dutra et al. (2014). RMFTs are phenomenological, microphysical theories of dense matter based on meson-exchange Lagrangians. We refer to a specific Lagrangian function as a single RMFT. By specifying a set of interaction constants within the RMFT, we generate a particular RMFT model that can be used to compute the corresponding EoS.
@@
Once the functional form of the Lagrangian and the nucleon-meson and meson-meson coupling constants are determined through fits to, e.g., properties of finite nuclei, nuclear matter, and astrophysical observations, e.g. Refs. Fattoyev et al. (2010); Steiner et al. (2013); Alford et al. (2022), no further parameters are needed to obtain finite-temperature Alford et al. (2023) or out-of-equilibrium Alford et al. (2024a, b) properties. The EoS obtained from an RMFT model is also thermodynamically consistent and causal. An RMFT model thus allows us to relate the microscopic properties of dense matter to astrophysical observables of neutron stars such as their masses and radii.
@@
We employ the methodology introduced in Essick et al. (2020b, 2021a, 2021b) as a way to avoid such biases and infer the breakdown scale with astrophysical observations. Although we focus on a specific RMFT in this paper, this method is generic and can be applied to any nuclear framework that is expected to become less reliable with increasing density. The key element is the flexibility to choose how model information from the RMFT EoS is used in constructing an EoS prior with which to analyze the data. We achieve this by incorporating information from an RMFT EoS into an otherwise model-agnostic approach for representing the EoS based on Gaussian Processes (GPs). The means and covariances of the GP are modified such that they closely follow RMFT EoSs generated using different choices of model parameters up to some value for the pressure. That pressure corresponds to a choice of a scale up to which we trust the RMFT to describe the properties of dense matter. Beyond that pressure, the GP smoothly transitions from an RMFT-informed kernel to a model-agnostic one. We construct multiple GPs for different choices of this transition pressure. The specific RMFT EoSs we use are fit to $\chi$EFT calculations. Therefore, when we incorporate nuclear theory information, the assumptions we make are strictly stronger than those of Refs. Essick et al. (2020b, 2021a, 2021b).
If the true EoS can be described by an RMFT at certain densities but not up to arbitrarily high densities, then we expect this to be discernible using astrophysical observations of neutron stars. We consider gravitational-wave observations of masses and tidal deformabilities Abbott et al. (2018, 2020), X-ray observations of masses and radii Miller et al. (2019a); Riley et al. (2019); Miller et al. (2021); Riley et al. (2021); Choudhury et al. (2024), and pulsar timing observations of heavy stars Antoniadis et al. (2013); Cromartie et al. (2019); Fonseca et al. (2021). We use a hierarchical inference scheme to infer the EoS under a variety of hypotheses about the breakdown scale of the RMFT framework. We consider both simulated and real astrophysical data, and study the extent to which we can infer the breakdown scale from astrophysical data alone. We find that with current astrophysical constraints, there is no strong evidence in favor of a breakdown of the studied RMFT. Nonetheless, we also find that a breakdown can be approximately identified using observations at around 3×3$\times$3 × the currently available measurement precision if such a breakdown is associated with, e.g., a strong, first-order phase transition.
@@ II.1 The RMFT EoS set
Commonly, an RMFT is further constrained by astrophysical observables. However, we do not want to make such an assumption a priori, as we would potentially use observations of matter that might not be comprised of nucleons at all. Thus, we do not require the RMFT EoSs to support $\sim 2\leavevmode\nobreak\ M_{\odot}$ stars. Since in our hybrid approach the RMFT description only holds at low densities, the resulting hybrid EoSs might still be able to produce high mass NSs (see e.g. Ref. Brodie and Haber (2023)).
@@
We produce 1109 EoSs using the method described in Ref. Alford et al. (2022). All EoSs come from the same RMFT Lagrangian density with different nucleon-meson and meson-meson coupling constants. The variation in these seven couplings is due to a fit to the $\chi$EFT energy per nucleon uncertainty band in neutron matter. All EoSs are causal, consistent with the saturation properties of isospin-symmetric nuclear matter, and consistent with $\chi$EFT at next-to-next-to leading order (N2LO) up to $1.5\,n_{0}$ Tews et al. (2018). Of the 1109 EoSs, only 90 predict neutron stars with $M\geq 2\,\operatorname{M}_{\odot}$. This set serves as a fiducial representation of RMFT EoSs. We will occasionally refer to the RMFT EoS set as a “prior”, by which we mean taking all EoSs from the set to be equally likely. More details are provided in Appendix A.
@@ II.3 Constructing an RMFT-informed GP prior and hybridization with a model-agnostic GP
We construct EoS priors conditioned on the RMFT EoS distribution up to several maximum (or “transition”) pressures as was done in Refs. Essick et al. (2020b, 2021a, 2021b) for $\chi$EFT. However, unlike $\chi$EFT, the RMFT is not automatically equipped with theoretical uncertainty estimates. That is, the distribution of RMFT EoSs introduced in Sec. II.1 does not include any estimate of systematic uncertainty from the fact that the RMFT Lagrangian includes a subset of all the possible terms. In contrast, $\chi$EFT systematic uncertainties are constructed from estimates of the truncation error introduced by only retaining terms up to a certain order in the EFT expansion. Therefore, at densities greater than saturation, the RMFT EoS distribution depends on the strategy we use to generate coupling constants from fitting inferred experimental results and ab initio calculations; see Sec. II.1.
@@
Equipped with a GP emulator for the RMFT EoS distribution, we then construct priors that closely follow the RMFT EoS distribution at low pressures but transition to more flexible model-agnostic priors above a transition pressure. References Essick et al. (2020b, 2021a, 2021b) achieved a similar effect by conditioning a model-agnostic prior on a $\chi$EFT emulator as if the emulator was an additional (noisy) observation. Nevertheless, we modify the approach here. Instead of conditioning an agnostic GP on the RMFT EoS emulator as if it was an additional observation, we instead construct a GP that exactly follows the RMFT EoS emulator at low pressures and then condition the agnostic GP at higher pressures. The new procedure guarantees that the resulting GP will only support smooth functions at low densities if the RMFT EoS emulator only supports smooth functions at low densities. It can be shown that the two approaches are equivalent if the (co)variance in the RMFT EoS emulator is very small compared to the (co)variance in the agnostic GP. See Appendix B for details.
@@
$E_{\rm SNM}(n)=E_{\rm bind}+K_{0}\left(\frac{n-n_{0}}{3n_{0}}\right)^{2}\,.$ (4)
We sample $E_{\rm bind}$, $K_{0}$, and $n_{0}$ from Gaussian distributions with mean and variance fit from the RMFT EoS set and given in Table 2. The nuclear energy per particle is equal to the total energy minus the contribution from electrons, which is known given the baryon number density and proton fraction
@@ III.1 Simulated Data
For X-ray data, we simulate 4 sources. We assume that uncertainties are uncorrelated and Gaussian on the mass and radius, which is reasonable for sources with independently measured masses from radio timing, such as J0740+6620 Miller et al. (2021); Riley et al. (2021) and J0437-4715 Choudhury et al. (2024). This assumption further improves as more photon counts and better background estimates are included,  Dittmann et al. (2024); Salmi et al. (2024), indicating that for future observations nearly-Gaussian uncertainties are plausible. We sample the uncertainties on mass and radius from a uniform distribution. The uncertainties in the most optimistic cases are $\sim 2$-3×3$\times$3 × smaller than current measurements. See Table 4 for the the simulated X-ray sources for the EoS consistent with the RMFT and Table 5 for the EoS with a phase transition.
@@ III.2.1 Recovering an RMFT EoS
Moving on to the posteriors, for all transition pressures, the simulated EoS (light blue) is recovered correctly, with an uncertainty that increases for lower transition pressures. This is expected because priors with lower transition pressure incorporate less information from the underlying RMFT and its strong model constraints. This effect is most severe where the simulated astrophysical observations are least informative, for example, at very high and low neutron star masses. The uncertainty in radius of a $2\,M_{\odot}$, for instance, is $\sim 1.25\leavevmode\nobreak\ \rm{km}$ with the model agnostic prior at $90\%$ credibility.
@@
For each transition pressure, $p_{\rm t}$, we compute the evidence $\mathcal{Z}(\{d_{i}\}|\mathcal{M}=p_{\rm t})$. This is to say, each GP conditioned at a different transition pressure forms a model $\mathcal{M}$, and we take a transition pressure of $p/c^{2}=10^{11}\rm{g}/\rm{cm}^{3}$ as a fiducial “agnostic value” since this incorporates the least information from the underlying RMFT. We then compute each model’s Bayes factor relative to the agnostic model and plot them in Fig. 3. There is a general preference for transitioning from RMFT to the more flexible GP at higher pressure, but it is weak with Bayes factors of $2$-$4$. Under the RMFT EoS prior, the Bayes factor relative to the agnostic model is higher, $\sim 11$. This is a consequence of our choice to condition the Bayes factors on the existence of heavy pulsars. EoSs with $M_{\rm TOV}\gtrsim 2.0$ effectively all appear in the posterior. Therefore, conditioned on the existence of heavy pulsars, the RMFT is more highly preferred than any agnostic model. In contrast, at $p_{\rm t}/c^{2}=3\times 10^{13}\,\rm{g}/\rm{cm}^{3}$ there is a dip in the evidence. This is because at this transition pressure, EoSs with low values of $R_{1.4}\sim 11.5\,\rm{km}$, which are inconsistent with the radius of the injected EoS ($R_{1.4}\sim 12.5\,\rm{km}$) are able to stiffen to reach $M_{\rm TOV}\gtrsim 2.1\,M_{\odot}$. Therefore low-radius EoSs are common even after conditioning on the existence of massive pulsars in this case, and their marginal likelihood is lower. If we remove this conditioning, then the RMFT is no longer substantially preferred relative to the agnostic model, and conditioning at $p_{t}/c^{2}=10^{14}\,\rm{g}/\rm{cm}^{3}$ is actually disfavored relative to the agnostic model despite the fact that the simulation EoS is itself an RMFT EoS.
@@
Figure 3: The Bayes factor of hybrid RMFT-agnostic GPs with various transition pressures compared to the “agnostic model” with a transition pressure of $10^{11}\rm{g}/\rm{cm}^{3}$. Points show Monte-Carlo estimates, with error bars showing ±$\pm 1$-$\sigma$ error in the estimate from the limited sample size of the EoSs. The 1-$\sigma$ region for the Bayes factor of the RMFT prior itself relative to the agnostic model is shown as a gray bar. There is an overall trend toward higher transition densities, but there is no conclusive preference for any transition pressure. The simulated data are consistent with all hybrid priors.
@@ III.2.2 Inferring a phase-transition EoS
The phase-transition EoS represents a point of low prior probability in all EoS priors. This leads to poorer sampling resolution than in the case of recovering an EoS without a phase transition, particularly for the maximally agnostic cases. [^2] See also Appendix C of Ref. Essick et al. (2023). For example, for the case of $p_{\rm t}=10^{13}\,\rm{g}/\rm{cm}^{3}$ and $3\times 10^{13}\,\rm{g}/\rm{cm}^{3}$ we have $\sim 20$ and $\sim 46$ effective samples, [^3] We define the number of effective samples $n_{\rm eff}\equiv\exp(S)$, where $S$ is the entropy of the underlying probability distribution $p(x)$, $S\equiv-\sum_{i}p(x_{i})\ln p(x_{i})$. respectively. Despite this, macroscopic observables are well recovered. In the cases above, $R_{1.4}=11.76^{+0.09}_{-0.11}\,\rm{km}$, and $R_{1.4}=11.73^{+0.12}_{-0.14}\,\rm{km}$ for $p_{\rm t}=10^{13}\,\rm{g}/\rm{cm}^{3}$ and $3\times 10^{13}\,\rm{g}/\rm{cm}^{3}$, respectively (quoted at $90\%$ credibility), consistent with the simulated value $R_{1.4}=11.72\leavevmode\nobreak\ \rm{km}$. For the case of a transition pressure $p/c^{2}=10^{14}\rm{g}/\rm{cm}^{3}$, however, the EoS is no longer recovered in that the posterior no longer reflects the simulation EoS within uncertainties. The maximum mass of all EoSs in the posterior is below the simulation EoS maximum mass. Furthermore, the posterior has only $\sim 2$ effective samples.

### v2

@@ I Introduction
Uncovering the equation of state (EoS) of dense matter is central to understanding neutron stars, bridging nuclear physics and astrophysics Oppenheimer and Volkoff (1939); Rhoades and Ruffini (1974); Lattimer and Prakash (2001); Lovato et al. (2022); Chatziioannou et al. (2024). While the EoS is in principle determined entirely by the standard model, in practice, there is currently no perturbative or exact approach to compute the properties of matter in the densest regions of neutron star cores. Therefore, for the purposes of interpreting macroscopic astronomical observables such as neutron star masses and radii, “model-agnostic” representations of the EoS minimize phenomenological model-dependence Landry and Essick (2019); Essick et al. (2020a); Legred et al. (2022). However, fully-agnostic approaches can overestimate uncertainties in regions where calculations from models of nuclear matter are reliable, for example, near nuclear saturation density ($n_{0}$), where a description of matter with nucleon degrees of freedom is applicable. In addition, extracting detailed microphysical information from model-agnostic approaches is not straightforward, because it is not usually clear what exact microscopic interaction is responsible for determining the EoS at different densities. Certain quantities are not accessible at all, such as the single-particle energies required to compute transport properties Schmitt and Shternin (2018). In contrast with model-agnostic strategies, a variety of phenomenological approaches have been used to model dense matter up to very high densities, but these come with varying degrees of (often difficult to quantify) systematic uncertainty. These can be broken down into two categories: (1) microscopic approaches such as Nambu-Jona-Lasinio models Nambu and Jona-Lasinio (1961); Buballa (2005) Skyrme models Skyrme (1961); Bender et al. (2003), relativistic mean-field theories Walecka (1974); Dutra et al. (2014), and parity-doublet models Detar and Kunihiro (1989); Jido et al. (2001), where systematic uncertainty arises from our limited knowledge of the relevant particles and their interactions, and (2) parametric approaches such as piecewise-polytrope models Read et al. (2009) and spectral models Lindblom (2010) that do not explicitly model microphysics and enforce a fixed functional form which may not reflect the true EoS Legred et al. (2022).
[…]
@@
Once the functional form of the Lagrangian and the nucleon-meson and meson-meson coupling constants are determined through fits to, e.g., properties of finite nuclei, nuclear matter, and astrophysical observations, e.g. Refs. Fattoyev et al. (2010); Steiner et al. (2013); Alford et al. (2022), no further parameters are needed to obtain finite-temperature Alford et al. (2023) or out-of-equilibrium Alford et al. (2024a, b) properties. The EoS obtained from an RMFT model is also thermodynamically consistent and causal because all thermodynamic quantities are consistently derived from a partition function obtained from the relativistic RMFT Lagrangian. An RMFT model thus allows us to relate the microscopic properties of dense matter to astrophysical observables of neutron stars, such as their masses and radii.
@@
We employ the methodology introduced in Essick et al. (2020b, 2021a, 2021b) as a way to avoid such biases and infer the breakdown scale with astrophysical observations. Although we focus on a specific RMFT in this paper, this method is generic and can be applied to any nuclear framework that is expected to become less reliable with increasing density. The key element is the flexibility to choose how model information from the RMFT EoS is used in constructing an EoS prior with which to analyze the data. We achieve this by incorporating information from an RMFT EoS into an otherwise model-agnostic approach for representing the EoS based on Gaussian Processes (GPs). The means and covariances of the GP are modified such that they closely follow RMFT EoSs generated using different choices of model parameters up to some value for the pressure. That pressure corresponds to a choice of a scale up to which we trust the RMFT to describe the properties of dense matter. Beyond that pressure, the GP smoothly transitions from an RMFT-informed kernel to a model-agnostic one. We construct multiple GPs for different choices of this transition pressure.
If the true EoS can be described by an RMFT at certain densities but not up to arbitrarily high densities, then we expect this to be discernible using astrophysical observations of neutron stars. We consider gravitational-wave observations of masses and tidal deformabilities Abbott et al. (2018, 2020), X-ray observations of masses and radii Miller et al. (2019a); Riley et al. (2019); Miller et al. (2021); Riley et al. (2021); Choudhury et al. (2024), and pulsar timing observations of heavy stars Antoniadis et al. (2013); Cromartie et al. (2019); Fonseca et al. (2021). We use a hierarchical inference scheme to infer the EoS under a variety of hypotheses about the breakdown scale of the RMFT framework. We consider both simulated and real astrophysical data, and study the extent to which we can infer the breakdown scale from astrophysical data alone. We find that with current astrophysical constraints, there is no strong evidence in favor of a breakdown of the studied RMFT. Nonetheless, we also find that a breakdown can be approximately identified using observations at around $3\times$ the currently available measurement precision if such a breakdown is associated with, e.g., a strong, first-order phase transition.
@@ II.1 The RMFT EoS set
Commonly, an RMFT is further constrained by astrophysical observables. However, we do not want to make such an assumption a priori, as we would potentially use observations of matter that might not be comprised of nucleons at all. Thus, we do not require the RMFT EoSs to support $\sim 2\penalty 10000\ M_{\odot}$ stars. Since in our hybrid approach the RMFT description only holds at low densities, the resulting hybrid EoSs might still be able to produce high mass NSs (see e.g. Ref. Brodie and Haber (2023)).
@@
[…]
@@ II.3 Constructing an RMFT-informed GP prior and hybridization with a model-agnostic GP
[…]
@@
[…]
@@
$E_{\rm SNM}(n)=E_{\rm bind}+\frac{1}{2}\kappa\left(\frac{n-n_{0}}{3n_{0}}\right)^{2}\,.$ (4)
We sample $E_{\rm bind}$, $\kappa$, and $n_{0}$ from Gaussian distributions with mean and variance fit from the RMFT EoS set and given in Table 2. The nuclear energy per particle is equal to the total energy minus the contribution from electrons, which is known given the baryon number density and proton fraction
@@ III.1 Simulated Data
For X-ray data, we simulate 4 sources. We assume that uncertainties are uncorrelated and Gaussian on the mass and radius, which is reasonable for sources with independently measured masses from radio timing, such as J0740+6620 Miller et al. (2021); Riley et al. (2021) and J0437-4715 Choudhury et al. (2024). This assumption further improves as more photon counts and better background estimates are included, Dittmann et al. (2024); Salmi et al. (2024), indicating that for future observations nearly-Gaussian uncertainties are plausible. We sample the uncertainties on mass and radius from uniform distributions of width $3\,M_{\odot}$ and $1.0\,\rm{km}$ respectively. The uncertainties in the most optimistic cases are $\sim 2$-$3\times$ smaller than current measurements. See Table 4 for the the simulated X-ray sources for the EoS consistent with the RMFT and Table 5 for the EoS with a phase transition.
@@ III.2.1 Recovering an RMFT EoS
Moving on to the posteriors, for all transition pressures, the simulated EoS (light blue) is recovered correctly, with an uncertainty that increases for lower transition pressures. This is expected because priors with lower transition pressure incorporate less information from the underlying RMFT and its strong model constraints. This effect is most severe where the simulated astrophysical observations are least informative, for example, at very high and low neutron star masses. The uncertainty in radius of a $2\,M_{\odot}$, for instance, is $\sim 1.25\penalty 10000\ \rm{km}$ with the model agnostic prior at $90\%$ credibility.
@@
For each transition pressure, $p_{\rm t}$, we compute the evidence $\mathcal{Z}(\{d_{i}\}|\mathcal{M}=p_{\rm t})$. This is to say, each GP conditioned at a different transition pressure forms a model $\mathcal{M}$, and we take a transition pressure of $p/c^{2}=10^{11}\rm{g}/\rm{cm}^{3}$ as a fiducial “agnostic value” since this incorporates the least information from the underlying RMFT. We then compute each model’s Bayes factor relative to the agnostic model and plot them in Fig. 3. There is a general preference for transitioning from RMFT to the more flexible GP at higher pressure, but it is weak with Bayes factors of 22-44. Under the RMFT EoS prior, the Bayes factor relative to the agnostic model is higher, $\sim 11$. This is a consequence of our choice to condition the Bayes factors on the existence of heavy pulsars. EoSs with $M_{\rm TOV}\gtrsim 2.0$ effectively all appear in the posterior. Therefore, conditioned on the existence of heavy pulsars, the RMFT is more highly preferred than any agnostic model. In contrast, at $p_{\rm t}/c^{2}=3\times 10^{13}\,\rm{g}/\rm{cm}^{3}$ there is a dip in the evidence. This is because at this transition pressure, EoSs with low values of $R_{1.4}\sim 11.5\,\rm{km}$, which are inconsistent with the radius of the injected EoS ($R_{1.4}\sim 12.5\,\rm{km}$) are able to stiffen to reach $M_{\rm TOV}\gtrsim 2.1\,M_{\odot}$. Therefore low-radius EoSs are common even after conditioning on the existence of massive pulsars in this case, and their marginal likelihood is lower. If we remove this conditioning, then the RMFT is no longer substantially preferred relative to the agnostic model, and conditioning at $p_{t}/c^{2}=10^{14}\,\rm{g}/\rm{cm}^{3}$ is actually disfavored relative to the agnostic model despite the fact that the simulation EoS is itself an RMFT EoS.
@@
Figure 3: The Bayes factor of hybrid RMFT-agnostic GPs with various transition pressures compared to the “agnostic model” with a transition pressure of $10^{11}\rm{g}/\rm{cm}^{3}$. Points show Monte-Carlo estimates, with error bars showing $\pm 1$-$\sigma$ error in the estimate from the limited sample size of the EoSs. The 1-$\sigma$ region for the Bayes factor of the RMFT prior itself relative to the agnostic model is shown as a gray bar. There is an overall trend toward higher transition densities, but there is no conclusive preference for any transition pressure. The simulated data are consistent with all hybrid priors.
@@ III.2.2 Inferring a phase-transition EoS
The phase-transition EoS represents a point of low prior probability in all EoS priors. This leads to poorer sampling resolution than in the case of recovering an EoS without a phase transition, particularly for the maximally agnostic cases. [^2] See also Appendix C of Ref. Essick et al. (2023). For example, for the case of $p_{\rm t}=10^{13}\,\rm{g}/\rm{cm}^{3}$ and $3\times 10^{13}\,\rm{g}/\rm{cm}^{3}$ we have $\sim 11$ and $\sim 33$ effective samples, [^3] We define the number of effective samples $n_{\rm eff}\equiv(\sum_{i}w_{i})^{2}/(\sum_{i}w_{i}^{2})$. Because the prior is uniform, this is closely related to the Kullback-Liebler (KL) divergence Kullback and Leibler (1951), $\mathrm{KL}(p||q)\equiv\sum_{x}p(x)\ln(p(x)/q(x))$. The KL divergence from prior to posterior for the PT-injection case recovered with the model transitioning at $1\times 10^{12}\,\rm{g}/\rm{cm^{3}}$ is $\sim 7.5$, the analogous KL value for the case with the RMFT injection is 3.7. This is another way of stating that the injected PT EoS is a poor representative of the entire EoS prior, since the posterior centered on it is far from the EoS prior. respectively. Despite this, macroscopic observables are well recovered. In the cases above, $R_{1.4}=11.76^{+0.09}_{-0.11}\,\rm{km}$, and $R_{1.4}=11.73^{+0.12}_{-0.14}\,\rm{km}$ for $p_{\rm t}=10^{13}\,\rm{g}/\rm{cm}^{3}$ and $3\times 10^{13}\,\rm{g}/\rm{cm}^{3}$, respectively (quoted at $90\%$ credibility), consistent with the simulated value $R_{1.4}=11.72\penalty 10000\ \rm{km}$. For the case of a transition pressure $p/c^{2}=10^{14}\rm{g}/\rm{cm}^{3}$, however, the EoS is no longer recovered in that the posterior no longer reflects the simulation EoS within uncertainties. The maximum mass of all EoSs in the posterior is below the simulation EoS maximum mass. Furthermore, the posterior has only $\sim 2$ effective samples.
@@
The cases of two possible EoSs with and without a phase transition represent two possible examples of astrophysically realistic EoSs. If the astrophysical EoS has a stronger phase transition than the one we simulated, we expect to find evidence with fewer data. We have chosen the parameters of the quark matter EoS in such a way that no unstable or twin branches arise, but a connected stable hybrid branch (case (c) in Fig. 2 of Ref. Alford and Han (2016)). This presents the greatest challenge, as twin or unstable branches would be easier to distinguish. If, on the other hand, there is a weaker phase transition, much more precise observations may be necessary to achieve similar evidence. To give a sense of the sensitivity of our analysis to the choice of the EoS phase transition parameters, we perform the same analysis with an EoS with a weaker transition in Appendix. D. We find that even in cases where astrophysical observations are unable to decisively identify a breakdown scale, they may still imply a posterior inconsistent with nuclear physics, in which case our approach will provide hints about the possible sources of the breakdown of the model.